### v1

# Modern approaches to building effective interpretable models of the property market using machine learning

Abstract
In this article, we review modern approaches to building interpretable models of property markets using machine learning on the base of mass valuation of property in the Primorye region, Russia. The researcher, lacking expertise in this topic, encounters numerous difficulties in the effort to build a good model. The main source of this is the huge difference between noisy real market data and ideal data which is very common in all types of tutorials on machine learning. This paper covers all stages of modeling: the collection of initial data, identification of outliers, the search and analysis of patterns in data, the formation and final choice of price factors, the building of the model, and the evaluation of its efficiency. For each stage, we highlight potential issues and describe sound methods for overcoming emerging difficulties on actual examples. We show that the combination of classical linear regression with interpolation methods of geostatistics allows to build an effective model for land parcels. For flats, when many objects are attributed to one spatial point the application of geostatistical methods is difficult. Therefore we suggest linear regression with automatic generation and selection of additional rules on the base of decision trees, so called the RuleFit method. Thus we show, that despite the strong restriction as the requirement of interpretability which is important in practical aspects, for example, legal matters, it is still possible to build effective models of real property markets.
keywords:
[…]
1 Introduction
The problem of forecasting the real estate value has existed for a long time and is actual nowadays. It could be approached differently and include different constraints that impact predicted values. For example, it could be the aim to get the most accurate predictions; it could be the limit on algorithm runtime; it could be the goal to get the interpretable model, and so on. The problems faced by researchers are the insufficiency of data; the market of interest is undeveloped; explanatory factors are insufficient for prediction of the real estate value and so on. There are also problems with the nonlinear influence of some variables on others or the low correlation between explanatory variables and dependent variable. All of the above complicate the prediction process and make it nontrivial.
The scope of research could be different – starting with the prediction of the price of one property unit up to a mass valuation of real estate in some geographic area. The last can be applied for research of the property market or administrative regulations such as property taxes. In most cases in the Russian Federation property taxes are determined from the cadastral value of the object which is calculated in the process of cadastral valuation. Cadastral valuation is a process of estimation of the cadastral value of property to create a clear and valid system of taxation and other relations governed by law. In the case of land parcels it could be a calculation of rental charge for land parcel use, the definition of land redemption value or buying from public domain or municipal property; for buildings and building spaces – calculation of taxes for buy/sell, rental and exchange of object, registration of succession rights, definition of exchange value, registration of gift agreement. The process of determination of cadastral value requires taking into account various factors, related to location, size, and quality of real estate property and also expert assessment of its potential for use in various purposes (Byrda, 2025). Cadastral valuation significantly influences the economic development of territories in municipalities. In the face of modern economic changes and financial difficulties, the importance of correct regulation of cadastral valuation as one of the factors stimulating the growth of municipal territories becomes particularly important. The precise cadastral valuation promotes the creation of a true picture of real market property values and leads to the improvement of the investment environment and stimulation of the outlays. Incorrect cadastral valuation or ineffective using results of cadastral valuation could reduce the investment attractiveness and give rise to a growth rate reduction of local economies and a decline of the social and economic situation. Overcharge of territories by unreasonably high values of cadastral value takes a toll on the interests of potential investors and may cause secular stagnation in the property market. Non-optimal instruments of cadastral valuation ignoring the physical-geographic factor of pricing instigates maldistribution of tax burden and as a result, the insufficient fundraising for covering municipalities’ needs. Real property tax based on valid cadastral valuation can be used for sponsorship of municipal projects focusing on the development of local amenities and the social sector. Incorrect cadastral valuation can prevent the development of particular areas and create financial and economic disbalance (Byrda, 2025).
Cadastral valuation is conducted by methods verging its predictions to market values as close as possible. The market value of an object can vary significantly over time. Due to this reason, the cadastral value of an object should be regularly updated through mass cadastral valuation of real estate. The result of such mass valuation is the model based on market data which predicts the price of the arbitrary object. In this paper, we discuss the methods of effective prediction for mass cadastral valuation of real estate on the example of state cadastral valuation. Initially, it means that such valuation should obey current law. In Russia, it states that the valuation process is regulated by methodological rules of mass valuation of real estate. According to them, the prediction of real estate value can be made using one of the next methods – comparative method, cost method, or income method. The comparative method is typically used in the case when the market of real estate property is highly developed and there is enough information about deals and offers. This method includes method of fiducial object and models of regression analysis such as linear regression and its derivatives (Ministry for Economic Development of Russia, 2018). However, in the modern world, where technologies improve every day and machine learning actively penetrates all spheres of life, making predictions more and more precise, it becomes clear that we need more widely implement such methods in the mass valuation of real estate to improve the existing algorithms.
Today there are many scientific works on the application of machine learning to prediction of property value. There are not only simple models such as linear regression and its variations (Adair et al., 1996) but also more complex ones, for example, Support Vector Machines (Phan, 2018) or a combination of several methods (Gu et al., 2011). In the review of mass valuation models (Jahanshiri et al., 2011), authors distinguish linear regression as a basic model and highlight its main limitation – it deals badly with spatial characteristics of data. It is important in the case of real estate property. For overcoming this problem authors suggest various spatial methods (Jahanshiri et al., 2011). But if models are created to form the base for taxing then transparency becomes the key aspect of the model. In this context transparency means interpretability. The results of the model should be presented in an accessible for ordinary person format – to explain clearly how input data influences on final results. In most cases, the machine learning models are black box – the programmer and users see the input data and the results but the mechanism of calculation is rather implicit. Appraiser can more deeply understand these calculation principles and algorithms, and what methods are used for the adjustment of estimates, but explaining it to an ordinary person who only wants to know the answer to just one question “How was the value of my property calculated?”, becomes almost impossible task. The interpretability of the model allows to find patterns and estimate how factors influence the price. It gives an opportunity to explain the constitutes of price to citizens who express disagreement with calculated cadastral values. Citizens can estimate the correctness of included factors and see their values for their property. In case of litigation, the presence of clear and sound cadastral values can help to protect it in law-court. Obviously, this argument is inapplicable in the case of black box models.
Nowadays there are many reference materials on machine learning, but most of them demonstrate concepts on the base of ideal data. These data significantly differ from real data that is subject to errors and fluctuation. As a result, when a researcher tries to practice ideas and methods from a book, he is faced with difficulties – either methods don’t work at all or work incorrectly. For solving this problem a researcher should undertake careful preparatory work with data. Unfortunately, tutorials do not pay much attention to this problem.
Taking into account all of the above this paper is written in an effort to create a guide on building a simple interpretable and effective model on real market data. In order to improve the predictive quality of the model we decide to extend the model with more complex concepts of machine learning – for example adding the rules influencing results, or to use the methods from a related area of science – geostatistics. Geostatistics provides a broad range of statistical models and tools for the analysis and processing of spatially distributed data (Cressie, 2015). Geostatistics finds an application in solving ecological, climatic, epidemiological, geological problems, and so on. However, principles of spatial modeling can be used for forecasting real estate value (Chica-Olmo, 2007; Demidova, 2013). We find the approach of modeling the general trend in the data with linear regression to be perspective. It allows to keep the interpretability of the model. The residuals of the model can be modeled by the kriging method which increases the precision of the model. Such an approach is known under the name regression-kriging (Hengl et al., 2007; Chica-Olmo and Cano-Guervos, 2020) and it is applied in the current work.
Also in this paper we present approaches for solving problems involving the identification of outliers, and the creation and selection of explanatory variables. Of note, we work with two market segments – flats and land parcels. For each segment, we analyzed the data and built the model. Sections 3 “Methods of identification of the outliers” and 4 “The formation and selection of explanatory variables” present the basic principles and interesting facts of the research performed for both segments. The section 5 “Modeling” consists of two parts describing each market segment. This section presents a detailed description of the methods and principles of modeling.
2 Description of target of research
The modeling territorially covers Primorsky Krai – the administrative region of the Russian Federation located in the southeast of the country (Figure 1). The region borders China on the west, North Korea on the south-west, and Khabarovsk Krai (Russian Federation) on the north. Primorsky Krai is the part of Far Eastern Federal District. The area of the region is 164 673 km2 which is comparable to areas of such countries as Bulgaria (110 994 km2), Greece (131 957 km2) or Cambodia (181 035 km2). The population of the region is 1 806 393, the administrative capital is the port city Vladivostok (population 591 628). Vladivostok is the political, cultural, scientific, and educational center of the region. In 2012 Vladivostok was the venue for the APEC (Asia-Pacific Economic Cooperation) summit. For this event a wide range of significant measures was taken – the development of telecommunication, the construction of new roads, enhancing the cultural and prestigious status of the region (building of Opera and Ballet Theatre, building of multipurpose health center), the reconstruction of energy and heat supply facilities contributing to the reduction of energy costs of enterprises and energy losses in public utilities. Vladivostok is one of ten cities in the country where the Federal University was created (its new campus on Russky Island was the venue for APEC). This university has become a powerful intellectual base for the region.
The membership of Russia in APEC and Primorsky Krai in particular as the main territory for interaction with other countries of the Asia-Pacific region allows balanced management of the regional economy and effective use of the natural resource potential of the Far East of Russia. This in turn increases the attractiveness of the region as a place for residence and creates opportunities for enhancing the business availability and for recruitment of potential investors (Chebotarev, 2012). Additionally, Asia-Pacific Economic Cooperation favors regional economic integration, supports a multilateral trading system, unlocks the potential of the digital economy, and helps to bridge the digital divides (Ministry for Economic Development of Russia, ).
All of the aforementioned indicate the development of Primorsky Krai and the economic growth of the region. Each sector of economics should work effectively. Because of this reason, the results of cadastral valuation must show the real market situation, create a valid taxation system, and be a base for other relations governed by law.
Precisely because of its favorable territorial location Vladivostok been in a third ten of Russia’s cities by population and separated by thousands of kilometers from the central part of Russia stably occupies a place in the top ten of cities of Russia by real property value. But the modeling of real property value is a difficult task due to several factors: complex terrain, historico-geographical urban development, various transport accessibility of different parts of the city, the existence of insular territories, the mixing of historic and modern buildings, nonuniform density of social, cultural and entertainment centers, coexistence of industrial estate and domestic development and so on. The requirement of interpretability of the model additionally complicates the task. Thus one should at each stage use the wide set of machine learning methods for building an effective model of property values for Vladivostok.
Machine learning methods are applicable in the case of enough data. They help to find the patterns and rules and to extend them to a general sample. Because of this reason, we consider only two mass market segments for building the models – flats and land parcels for private housing projects. The object for predictions is the real estate value, specifically – the Per Square Meter Price of real estate value (PSMP):
Basic data for the real estate value were taken from open sources (like websites with offers for the sale of property) and deals. Information about deals was presented by the Ministry of Property and Land Relations of Russia according to federal law. We use data collected for one year before the modeling. Thus we formed two samples for modeling – 26432 objects for segment “Flats” and 19246 objects for segment “Land parcels”. For each segment, we create and analyze explanatory variables and their correlation relationships, search the patterns, and build models.
3 Identification of outliers
An important step in modeling is checking data for validity. The data is prone to mistakes of several origins. The most common of them are accidental mistakes (typos and so on) and intentional distortion of data by publishing fictitious values – too high or too low. As a result, the process of creating a model becomes more difficult and may lead to incorrect results. It’s worth noting that in the process of analyzing the data on the outliers, one should take into account the type of objects. A land parcel is basically a one-dimensional object. Usually, the price of such an object depends on its distance relative to other objects, for example center of the town. On the other hand, the value of a flat depends not only on its location but also on the characteristics of the multi-unit residential building where the flat is located. In the second case, the process of identification of outliers can be more difficult. In the case when we analyze the one-dimensional objects (land parcels), we can plot their per square meter price on a map and estimate their validity from general considerations.
Figure 2 shows that the points located close to each other can have significantly different per square meter prices. This may be caused by individual characteristics of the object – the land parcel could be in a zone with a high risk of flooding or the land parcel could have a significant slope. All these factors make land parcel exploitation more difficult. However, the determination of the cadastral value of the object is the mass process reflecting the value of a typical object without delving too much into its specific characteristics. Because of this, reason we can expect a similar per square meter price among land parcels in some local area that changes with distances to other objects included in price factors (center of a town, water objects, social objects, etc.). For the exclusion of outliers, we can start with the clusterization of points on the base of their spatial relationship. The next thing we can calculate the median or average value for each cluster and find the outliers using such methods as box-plot or z-score or so on. Also, such clustering helps to take into account the inhomogeneity of administrative territory. For example, the territory could include the city and the suburbs where a per square meter price is sufficiently different (Figure 3). On the contrary, if we consider the whole territory without clusterization we can lose part of the data because firstly the per square meter price of the city is much higher than the per square meter price of the suburbs and secondly the number of data points in the city are much less than the corresponding number in suburbs. In this case, the land parcels located in the city would be outliers and by excluding them we will have major losses of important data and our predicted value would be sufficiently distorted.
Also, we can find the outliers using the probability density function of dependent variable. For example, Figure 4 shows that values between 50000 rub and 100000 rub for the deals are doubtful.
We can confirm this assumption by taking buildings and analyzing per square meter prices in each building. The example of one building is shown in Table 1 where suspicious value is highlighted. Such deals could be fictitious and if we wouldn’t exclude them they would have a negative effect on the forecasting.
Type | Storey of flat | Total price, rub | Area of flat, m2 | PSMP, rub |
offer | 5 | 3 850 000 | 17,1 | 225 146 |
offer | 2 | 4 500 000 | 23,1 | 194 805 |
offer | 8 | 6 150 000 | 34,5 | 178 261 |
offer | 9 | 4 350 000 | 17,0 | 255 882 |
offer | 3 | 4 200 000 | 17,0 | 247 059 |
offer | 4 | 3 800 000 | 16,7 | 227 545 |
offer | 5 | 3 800 000 | 17,0 | 223 529 |
deal | 5 | 4 300 000 | 16,7 | 257 485 |
deal | 2 | 4 100 000 | 23,0 | 178 261 |
deal | 4 | 4 050 000 | 16,9 | 239 645 |
deal | 4 | 900 000 | 16,8 | 53 571 |
deal | 4 | 4 900 000 | 23,3 | 210 300 |
deal | 4 | 3 800 000 | 16,7 | 227 545 |
More serious methods such as machine learning also can be used for the identification of outliers. Particularly it would help for many-dimensional objects e.g. flats. Firstly we should perform territorial clustering of points that results in the division of the populated place into regions (clusters). Next, we create the new clusters in each region on the base of the characteristics of the buildings e.g. “Wall material”, “Number of stories” and “Year built”. This procedure allows to highlight typical buildings. We can identify such clusters either manually on the basis of our expertise and knowledge of historic periods of city development or we can use machine learning methods implementing clustering algorithms. Such algorithms take as input the buildings in the space of the above-mentioned characteristics and return the class markers for each building. In this paper, we use the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996). This algorithm is based on the idea of closeness between points in the space of features. It means that DBSCAN groups the close points and marks as outliers the single points. The minimal number of points in a neighborhood for a point to be considered as a core point and the maximum distance between two points for one to be considered as in the neighborhood of the other are adjustable parameters of the algorithm (Moreira et al., 2005). By changing the first we can control the density of resulting clusters while the second parameter is the most important one and should be chosen accurately depending on the data and the metric. The metric specifies how the distances in the space of features are calculated. After the process of finding typical buildings is finished one can start the identification of outliers within each group. During the analysis, we noticed that there is a negative dependence between the per square meter price and the area of the flat. In addition, this dependence is not linear but rather exponential (Figure 5). The highlighting of the trend for such data is a nontrivial task. If there are signs of linear dependence then we can try RANdom Sample Consensus (RANSAC) regression (Kumar, 2020; Fischler and Bolles, 1981) – linear regression robust to outliers. This algorithm is an iterative procedure where every step assumes the forming of the sub-sample and fitting linear model. Then this model is applied to the total sample and residuals are calculated. Everything beyond the residual threshold is marked as outliers. Then the model retrains on other values and all steps are repeated. The new model and previous model are compared on the number of outliers. A model that contains a smaller number of outliers is preferable. In our case, we can transform the exponential dependence to linear by taking the natural logarithm of data. So for each cluster found by DBSCAN, we apply RANSAC regression to find the outliers and to highlight the trend. The dependent variable is the natural logarithm of the per square meter price and the explanatory variable is the area of flat. This procedure allows to find the trend for data with high fluctuations and noise and hinges on the reasonable assumption that flats in buildings with similar characteristics located within some compact territory should have similar per square meter prices. The points that are highly different from the trend are considered to be outliers. These values can be explained by the individual characteristics of the flat and can’t account for the mass valuation. The example of the results of RANSAC regression is shown in Figure 5, the points marked as outliers are not used for forecasting.
4 Formation and selection of explanatory variables
According to methodological rules of mass valuation of real estate, we can use for explanatory variables such information as general characteristics of an object (object’s purpose, area, object’s name, wall material, year build, extension of the land parcel, kind of soil and so on); information about a location of an object (a distance to the roads, to the water objects, to a recreation area, a city center and so on); information about engineering infrastructure (a class of linear object, a distance to a power line, a rated power of a linear object and so on); other matter (a price level of a commodity bundle, commodity circulation per 1 person, a distance to extractive industry territory) (Ministry for Economic Development of Russia, 2018).
General information about the object is contained in the Unified State Register of Real Estate and in accordance with the law is given by the ministry to the certified organization which makes evaluation. Socioeconomic data can be found in the annual reviews of the development of the territory. However such information can be useful only for large territories that contain a few municipal districts or a few towns. Geographical features can be extracted with the help of mapping software – NextGIS and Arcgis. Also, we can get such information using programming tools, for example, Python and open non-profit mapping source – OpenStreetMap. Using the library “osmnx” from Python we can get all kinds of socially significant facilities – nursery schools, schools, hospitals, shops, malls, and so on. Then we can calculate the straight-line distances to these objects or, for example, count the number of such objects in some radius. We can use this data as explanatory variables for a predictive model. Also libraries “osmnx” and “networkx” allow loading road network graphs. So we have the opportunity to calculate the distance by the roads. This calculation improves the accuracy of the variables. It can be important when we have a territory with irregular coastlines (Figure 6) or deal with islands. In the last case, we can use adjusting factors reflecting the accessibility of the island.
Also, we can aggregate some variables into one (as explained below in the discussion of the correlation matrix) or use conceptions from the related fields, for example, graph theory. Thus we created the factor “Development of the road network” that helped to estimate the location of object in relation to transport accessibility. If the object has many crossroads and roads nearby then this object is more attractive and therefore more expensive. This idea is more important for big cities. For the derivation of the factor “Development of the road network”, we used the road network graphs and the conception of centrality. The factor of centrality in graph theory and road network analysis allows to find the most important nodes of a graph (Newman, 2010). Important nodes are points that have the highest number of shortest routes. For the calculation of centrality, we consider the points of crossroads and the endpoints of roads as nodes of a graph and roads between these points as graph edges. In these terms the importance of a node means how many shortest roads between two points go through this node. When we calculate the shortest routes we naturally associate the road distance between two points with weights of the corresponding graph edges. We use the Python library “networkx” to calculate the centrality. This library allows to calculate the centrality in different ways. In this paper, we use harmonic centrality (Marchiori and Latora, 2000) because it is simple for calculations and understanding. The harmonic centrality of a node is the sum of inverse lengths of the shortest routes for each node to a particular node. It has a perfect intuitive interpretation – smaller distances give a higher contribution to centrality. Mathematically the harmonic centrality can be written as:
where - the node for which the centrality is calculated; - the other nodes of the graph; - the shortest route between and . The shortest route between two points is the route for which the sum of weights of constituting edges is minimal.
The roads can have bi-directional and unidirectional traffic. Because the graph of the road network is an oriented graph, it means that every edge of the graph has a direction. We should take this fact into account when calculating the centrality. We calculated the harmonic centrality for the road network of Vladivostok (Figure 7 (a)).
(a) | (b) |
The mapping between the centrality of nodes and the centrality of objects for evaluation could be done in several different ways ordered from the simplest to more sophisticated:
1.
first step: find the closest node of the road network to the object;
second step: consider this result as an estimation of the centrality of the object. -
2.
first step: create the buffer area of some radius around the object;
second step: calculate how much points of the road graph lie inside this area;
third step: calculate the mean value of the centrality among these points;
fourth step: consider this result as the estimation of the centrality of the object. -
3.
Also, we can do the extra calculations – interpolate over nodes of the road graph. Interpolation is the method of finding unknown values of some function in intermediate points using a discrete set of known values of a function at some points.
We implement the third approach because after the interpolation we get the smooth continuous surface of centrality values at any point. It allows to estimate the centrality of our objects more precisely compared to other mentioned possibilities. For interpolation, we use the method of basis functions. In this method, we get estimation as a linear combination of basis functions (Dem’yanov and Savel’eva, 2010; Buhmann, 2000; Regis and Shoemaker, 2006; Zhang et al., 2024).
where - estimation at the point ; - distance from point to point ; - basis function that depends on the distance; - weight coefficients. Coefficient defines algebraic sign of an element and its contribution.
We use the linear spline function as the basis function, which has the form:
The weight coefficients are derived from the condition that at known points the results of interpolation must coincide with known values, i.e. the interpolation model in all points must estimate values equal to . Thus we need to solve the system of linear equations with unknown variables.
To do this we use the function RBFInterpolator from Python library ”SciPy”. The results of interpolation for Vladivostok are shown in Figure 7 (b). Here for the centrality of the objects we use the term “Development of the road network” underlining our mathematical interpretation of this concept.
If we want to build a model for an inhomogeneous area (for example if territory contains a few towns or villages) then we can try to introduce new variables, using aggregation of existing variables. For example, we created the factor “Population of a residential place divided on the distance to Vladivostok by the roads”. On the one hand, the increasing of distance between the residential place and Vladivostok’s administration negatively affects the per square meter price; on the other hand, the population of the residential place positively affects the per square meter price. These considerations led us to create the factor that aggregates two variables – the population of a residential place and the distance to Vladivostok by the roads. Econometrically this factor should have a positive correlation with the per square meter price – with an increase in the population of residential place and a decrease inf the distance to Vladivostok the per square meter price increases. Figure 8(a) shows the relation between this factor and the average value of per square meter price over a building. We can notice that data contains positive a linear correlation however we also can notice that there are quite a lot of points when the aggregate factor has low values. Because of this, we tried to raise this factor to different powers. Figure 8(b) shows the relation between this factor raised to 1/3 power and the average value of per square meter price over a building. Such transformation gives a better value of correlation.
(a) | (b) |
The next important step in the way of creating the model is the estimation of correlations of all factors with the dependent variable (per square meter price) and the selection of the most important of them. For this step, we created the matrix of correlation for quantitative factors and we analyzed probability density functions of per square meter price for categorical factors.
When we study the matrices of correlation (Figure 9, Figure 10) we are interested not only in the relationship between explanatory variables and dependent variable but also in the relationships between explanatory variables with each other. The first model that we tried was the linear regression, because of its interpretability. The linear model has a set of assumptions, one of them is the absence of multicollinearity in data. Multicollinearity is a situation when two or more variables have a significant correlation with one another (). In this case, the inclusion of them in the model becomes unpractical. Because of this multicollinearity should be removed from data before modeling. We can achieve it in several ways: we can keep only factors that have a more significant correlation with the dependent variable; we can drop factors with the wrong sign of correlation with the dependent variable, which means that such factors have incorrect economic substance; finally, we can integrate several factors into one employing the principal component analysis (PCA) (Ajvazyan, 2010). This method is useful when several factors have similar economic substance. For example, the distance to school, nursery school, and supermarkets can be unified into one factor which means the distance to socially significant facilities (alternatively instead of distance we can use the number of objects). The reason for such aggregation is that these factors have significant relation with one another, but if we drop one of them we will lose some information. Instead, we incorporate them into the principal component which includes the dispersion of all these factors allowing to save more information about our variables. It is worth noting that the majority of considered features reflect spatial characteristics of data. The inclusion of such features in the model is critically important for building an effective model of spatial data (Jahanshiri et al., 2011) like land parcels and flats.
Previously we noted that before including categorical factors into a model we plot probability density functions of a per square meter price. A similar method of analysis would be drawing of a box-plot diagram. Traditionally the categorical factors are included in a model as binary features and one should pay attention not only to average values of per square meter price for each group but also to the representativity of the group in the data. For example, Figure 11 shows the probability density function for the factor “Wall material” and we see that flats in timber buildings have a per square meter price far less than the other materials. For this reason, we can’t include this binary feature in a model due to its nonrepresentativity. A possible solution to this problem can be the introduction of a correction factor reducing the per square meter price of such objects after getting of estimation of a model.
After completing all preparative steps – removing multicollinearity; dropping the factors without the economic substance; and dropping the non-significant features, we can move to the question of selection of price-forming factors that we may include in the model. There are different ways to solve this problem (Ajvazyan, 2010) – we can try all possible combinations of factors and choose the best model taking into account some requirements (minimizing errors or maximizing ); we can choose a model that meets the initial conditions for its using (for example for ordinary least squares such conditions are unbiasedness, consistency and effectiveness of estimations); also we can use algorithms for automatic feature selection – Forward selection or Backward elimination (Narisetty, 2020).
For segment “Land parcels” where the correlation of features with dependent variable has sufficiently high values (Figure 9) we applied Forward selection of features and also we took into account the value of and realization of the basic conditions for ordinary least squares method (Ajvazyan, 2010; Kutner, 2005) (Durbin–Watson statistic, Jarque-Bera test, F-test).
Number of factors | D–W stat | J-B (prob) | F-stat (prob) | |
1 | 0,636865 | 1,238017 | 0,000005 | 1,173483e-207 |
2 | 0,712070 | 1,371189 | 0,072956 | 3,926066e-253 |
3 | 0,749263 | 1,564183 | 0,175133 | 1,023501e-279 |
4 | 0,752834 | 1,564116 | 0,073893 | 2,622868e-281 |
5 | 0,763181 | 1,609770 | 0,002961 | 1,090438e-288 |
6 | 0,773200 | 1,566779 | 0,004079 | 3,415025e-296 |
From Table 2 one can see that in this particluar case, 4 factors are the optimal choice for modeling linear regression taking into account the conditions for using ordinary least squares.
5 Modeling
In this paper, we show the results of the prediction of the cadastral value of objects located in Primorsky Krai – a sufficiently large region of the Russian Federation. The area of this region is 165 thousand square kilometers. The region has a capital city – Vladivostok (with a population of about 600 thousand people), towns (with a population of not more than 150 thousand people), and villages. All these territories have different property markets, price levels, and quality of life, that depend on economic and territorial characteristics. The methodological rules (Ministry for Economic Development of Russia, 2018) recommend splitting the territory on the base of administrative boundaries of regions taking into account the population. This recommendation can be applied to objects that are located on homogeneous territory, for example for flats that are located in one city. However, this recommendation works poorly for objects that are spread over a large area e.g. land parcels. They occupy a large heterogeneous area – the city and the suburbs and often form allotment societies (also called gardeners’ non-commercial partnerships). Gardeners’ non-commercial partnerships are a collective form of gardening. The members perform their activities on an individual basis. Activities of such partnerships are regulated by the federal law. The partnerships can have their regulations, convoke meetings, and appoint a chairman. Similar organizations exist in many countries, for example, Germany, Sweden, Poland, and the USA. In every country, there are own procedures for the registration and obligation to pay a tax. In Russia, such partnerships must pay taxes, in particular land tax (Seregina, 2019) – the tax paid by physical persons or organizations that have property right of land plot or right of land plot perpetual usage or right of lifetime inheritable possession. Tax amount is defined from cadastral value and taxation rate. Such an allotment society can include different administrative regions. In such cases, it is better to use a grouping of objects employing clusterization algorithms instead of using formal administrative borders. Most of the algorithms are based on spatial relations between objects. Because of this reason, for the segment “Land parcels” we used a k-means clustering algorithm (MacQuenn, 1967), that takes into account the spatial positions of objects. This method is simple for understanding and for realization. It recognizes the accumulation of points and creates the clusters in such a way that the distance to the center of the cluster for every point must be minimal. By using this algorithm we are more confident that neighboring points belong to one cluster.
All further calculations of modeling were done for the capital region – Vladivostok and its suburbs.
segment “Land parcels”
First of all, we plotted the probability density function of the dependent variable (per square meter price) and noticed that it is not normal (Figure 12(a)). Although statistic and machine learning theories don’t have the strict rule requiring the normality of the probability density function of the dependent variable, usually analysts prefer to work with data distributed approximately normally. Because of this reason, we used a natural logarithm of the per square meter price (Figure 12(b)). Also such transformation of data guarantees that prediction of modeling will have only positive values.
(a) | (b) |
We start the modeling by trying the simple and interpretable model – the linear regression. As we mentioned above for the selection of features we apply Forward selection using a correlation matrix (Figure 9). The data was split into two parts – a training set (70% of data) and a testing set (30% of data). Figure 13 shows the result of the training of the model.
Using Figure 13 we can check the model on basic requirements for linear regression. We see that the coefficients of variables are significant (); in general, the equation of regression is significant (); the coefficients of variables have the right sign according to economic substance; model residuals have normal distribution () and don’t have autocorrelation (); adjusted coefficient of determination is significant () and it is not differ sufficiently from .
Also factors included in the model don’t have multicollinearity because the value of the variance inflation factor (VIF) is less than 10 for each factor (Shrestha, 2020) (Table 3).
Factors | |
Distance to the water objects (straight-line distance) | 1,463 |
Distance to the seacoast (straight-line distance) | 1,320 |
Distance to the town hall (straight-line distance) | 1,936 |
Aggregate: Distance to the hospitals (straight-line distance) & Distance to the closest nursery school (straight-line distance) & Distance to the closest school (straight-line distance) | 1,424 |
Figure 14 shows the distribution of model residuals. This distribution is approximately normal which satisfies the assumptions of linear regression. Also, we tested the normality of residuals using the library “SciPy”. This test compares calculated with the significance threshold (for this paper we chose ). If calculated is more than we can assume that the distribution of model residuals is normal (in our case ).
Also, we created the scatter plot values of residuals from their numerical order. From Figure 15 we can see that the data does not contain any dependencies. It means that residuals are homoscedastic and the other assumption of linear regression is satisfied.
On the base of the described results, we can state that our model is effective and interpretable. As we’ve already said the key factor for the identification of the price of a land parcel is its location. We see that linear regression does relatively well but the price of an object can include latent (local-territorial) characteristics. In order to take this into account we tried to apply a powerful instrument dealing with spatially distributed data – geostatistics. Geostatistical methods allow to model the distribution of objects in space. The central task of geostatistics is to reconstruct a studied phenomenon on the base of the values calculated in a limited number of points (Dem’yanov and Savel’eva, 2010). Mathematically we can contemplate this problem as an interpolation problem. For getting the best spatial estimations kriging methods are used. Estimations obtained by kriging have a minimal variation of error. Thus kriging is the best estimator in a class of linear interpolators (Dem’yanov and Savel’eva, 2010). There are two basic conditions for using these methods – data should have spatial continuity and stationarity (Kovalevskiy, 2012). The first condition means that two points located close together should have rather similar values than points separated more. The second condition means that in some locality the studied phenomenon must be spatially homogeneous.
We can check the condition of stationarity using the probability density function of a variable, in our case we analyze the distribution of residuals of our linear model. If the histogram is similar to the Gaussian distribution then in all likelihood there is stationarity (statistical homogeneity). Strictly speaking, this criteria is not a sufficient condition for confirmation of stationarity but it is strong enough and sufficient condition for practical work (Kovalevskiy, 2012). Figure 14 and the result of the Jarque–Bera test confirm the stationarity in linear model residuals. It’s worth noting that before checking data on stationarity we need to examine whether data contains a trend. In our case, we removed the trend from data using linear regression and then we modeled its residuals with geostatistical methods. The trend in data can distort the variogram – an instrument for estimation of spatial correlations in data (Dem’yanov and Savel’eva, 2010; Isaaks, 1989). The variogram model is a function describing a dependence between the studied spatial variable and the distance. If data contains a trend then a variogram would reproduce a large-scale trend that leads to missing a correlation of the observed variable on a small scale.
Also yet one important criteria (but also not sufficient) of stationarity is that the variogram should have a sill beginning at some distance. For building a variogram the researcher can choose the fitting curve from a list of basic models of variogram. In this paper, we used the exponential model, which is shown in Figure 16.
The variogram is used in finding the estimations of ordinary kriging (Dem’yanov and Savel’eva, 2010; Isaaks, 1989) and substituted instead of covariation function in the process of minimizing the Lagrangian that allows to get interpolation coefficients of ordinary kriging taking into account imposed on them constraints (Dem’yanov and Savel’eva, 2010).
The second condition for implementing kriging – the spatial continuity. We checked it by plotting the correlation of points as a function of the distance between them. From Figure 17 we see that with increasing distance between points, the correlation decreases. Therefore this criteria is satisfied.
Thus the final value of the per square meter price is the sum of the trend that we got using the linear regression and the prediction of linear model residuals that we got using kriging.
The last step of modeling is getting the resulting predicted estimations – we need to take the exponent of the predicted logarithmic value.
We assess the result by comparison of fitting errors using cross-validation. The average fitting error for linear regression is 28,36% and for linear regression with kriging is 19,23%. Also in Figure 18 predicted vs. actual plots are shown. Both clouds are arranged along the diagonal line but for regression-kriging it has a more elongated shape corresponding to lower errors.
(a) | (b) |
segment “Flats”
From the correlation matrix (Figure 10) we can see that correlations between the dependent variable and pricing variables for the segment “Flat” are significantly smaller than for the segment “Land parcels”. There is no explicit dependence on the object’s location. Nevertheless, we see that socially important factors such as “Number of shopping malls (within a radius of 1 km)” and “Number of cafes (within a radius of 1 km)” have higher correlations with the per square meter price than other factors, but still insufficient. However, we start with an effort to create a simple and interpretable model – linear regression. This step is necessary according to methodological rules of mass valuation of real estate (Ministry for Economic Development of Russia, 2018). Before modeling we did data preprocessing – we selected the factors (features) included in the model and made data standardization (Konayeva et al., 2024). The standardization is useful in the case of considerably different values of factors in order to reduce them to one scale. As we mentioned earlier the selection of the variables can be done in various ways according to theoretic assumptions and/or assumptions of area under the study. In this paper we used a machine algorithm of selection – SelectKBest (library “sklearn”). This method estimates the dependence of factors with a dependent variable using statistic tests and selects the best factors according to preferred criteria. We used the F-test as a statistical test for estimating the linear dependence between factors.
We selected 8 the most significant pricing variables using SelectKBest – “Area of flat”, “Distance to the water objects (straight-line distance)”, “Distance to the seacoast (straight-line distance)”, “Distance to the closest railway station (straight-line distance)”, “Distance to the area of negatively influencing factors (straight-line distance)”, “Number of cafes (within a radius of 1 km)”, “Distance to the town hall (distance by the roads)”, “Wall material: large-panel”. We created the linear regression model using these factors and taking as the dependent variable the per square meter price without any more transformations. The results of the model are shown in Figure 19.
From Figure 19 we can see that the coefficients of variables are significant (); in general the regression is significant (); the coefficients of variables have right sign according to economic substance; model residuals don’t have autocorrelation (); the value of Condition Number is 2,96, it means that factors don’t contain multicollinearity.
Also, important conditions for using a linear regression are expected value of residuals must be equal to zero (in our case we got the value 3,61e-10) and the dispersion of residuals must be constant (Figure 20) without any trend.
But the residuals of our model are not normally distributed () (Figure 21), which means that estimates got by the ordinary least squares are not optimal – they don’t have minimal dispersion in the class of linear unbiased estimators. Also in this model we have relatively low values of (), which means that the explanatory power of the model is poor.
Figure 22 (a) shows the comparison of predicted and real values of per square meter price. In an ideal case, the cluster of points is expected to be an ellipse stretched out along line y=x, which means that our predictive values should approach real values with the least dispersion. But from Figure 22 (a) we can see that in most cases our predictions overestimate the low values and underestimate the high values.
(a) | (b) |
We summarized the results of the linear regression model in Table 4. We can use it for comparison with other models.
train dataset | test dataset | train dataset | test dataset | Average approximation error (train dataset) | Average approximation error (test dataset) |
0,319 | 0,339 | 17 890,32 | 17 947,46 | 11,7% | 11,9% |
Sometimes the data can include non-linear dependencies that make a linear regression ineffective. In such a situation, we can try other algorithms that work better with non-linear dependencies in data. The most popular method is the “Decision tree” (Díaz-Ramírez et al., 2023). This algorithm forms the hierarchic tree structure that consists of rules “If … then …”. The rules are generated on the base of the training dataset in the process of training. The rules of the algorithm are created on natural language and they are simple for understanding. This is an advantage of the algorithm. But also method has significant shortage – it is prone to overfitting. It means that generated rules will fit only the training dataset, this tree will have bad results on other datasets. The extension of this algorithm called “Random Forest” (Breiman et al., 2017) rectifies this shortcoming. This method creates a set of decision trees. The advantage of it is the high accuracy of prediction and also this algorithm tends to work well with hyperparameters by default. Compared to the “Decision tree” the “Random Forest” is much less prone to overfitting. The shortage of this method is the complicated interpretation of modeling results due to the large number of generated rules.
Yet one algorithm that predicts with good accuracy and possesses better interpretability than “Random Forest” is “RuleFit” (Molnar, 2022; Masís, 2021). The underlying concept of the method is a combination of linear regression and predicted models that are based on decision trees. This model was developed in 2005 and was positioned as a sparse linear regression model (a class of models that after the training can have coefficients with zero weights). This algorithm detects the effects of interaction between features in the form of decision rules. RuleFit automatically generates these functions (rules) from decision trees. Using RuleFit we can choose any algorithm that generates the set of decision trees, for example, random forest or gradient boosting. Generated rules are used in the linear regression model as individual variables. The base model of the algorithm is LASSO (Least Absolute Shrinkage and Selection Operator also known as L1 regularization) (Tibshirani, 1996), because it allows the selection of the most significant rules from the set of rules for including them in the final model. LASSO is similar to ordinary linear regression but there is a supplementary component of regularization in the optimization functional of the model. This component is the penalty for the complexity of the model. Mathematically we can write this as:
where - Residual Sum of Squares; - regularization parameter; - coefficients if models. The goal of minimization is reaching a compromise between errors of regression and the dimension of space of features. Therefore RuleFit is a LASSO regression that takes as input both individual factors and the rules created by the random forest algorithm. These rules describe the interplay of variables. The distinctive characteristic of LASSO is going to zero of some factors as a result of minimization. This procedure determines the most informative factors. The tendency of some factors to disappear is controlled with a regularization parameter that is fitted using cross-validation (procedure of empiric estimation of the generalization ability of the algorithm). More information about adjusting the regularization parameter can be found here (Friedman et al., 2010). Due to the fact that the algorithm automatically selects the most significant factors the additional checking of selected factors is not required. Unlike the Ordinary Least Squares (OLS) method, the LASSO (and consequently RuleFit) does not have any prerequisites. We can estimate the effectiveness of the model by using common evaluation metrics such as (Mean Squared Error), (Mean Absolute Error), (Average approximation error), and (Coefficient of determination).
For building the RuleFit model we used both binary and continuous features. As continuous features, we used the factors from the correlation matrix (Figure 10) where correlation with the dependent variable is 0,15 and more. Factors satisfying this are:
: Area of flat
: Distance to the hospitals (straight-line distance)
: Distance to the water objects (straight-line distance)
: Distance to the seacoast (straight-line distance)
: Distance to the business activities center (straight-line distance)
: Distance to the closest railway station (straight-line distance)
: Distance to the area of negatively influencing factors (straight-line distance)
: Number of shopping malls (within a radius of 1 km)
: Number of cafes (within a radius of 1 km)
: Distance to the town hall (distance by the roads)
: Aggregate: Year built & Number of storeys
As binary features we used the factors describing the storey of the flat (“the first”, “the last”); the wall material of the building (“bricky”, “large-panel”, “monolithic”, “small-element concrete blocks”; at the same time the wall materials such as “made of timber” and “not determined” were excluded from consideration because they are nonrepresentational and don’t have economic substance). Thus we used 17 features for modeling. We created all kinds of combinations of these factors using RuleFit (we create models with the number of factors from 8 to 12). We created 86632 () models in total. For a selection of the best model we used values of and ; and also we checked the signs of variables according to their economic substance. In order to keep the interpretability of the model we applied a limit on the number of rules as a hyperparameter of the model – 50 rules. For a generation of rules we used the Random Forest algorithm. As a result of LASSO the part of rules were evaluated as insignificant and excluded, and only significant were kept. The final model contains 22 rules and 12 features. They are shown in the Table 6. Due to its volume, we present it in the A.
This model contains both the factors describing the basic characteristics of a building – aggregated variables of the year built and the number of storeys, wall materials – allowing to differentiate one building from another, and factors characterizing the location of building – “Distance to the seacoast”, “Distance to the closest railway station”, “Distance to the water objects”, “Number of cafes”, “Distance to the town hall”. Also, the model distinguishes the individual characteristics of a flat within the building including features such as “Area of flat”, “Storey of flat: the first”, and “Storey of flat: the last”. We can note that the majority of rules include factors “Distance to the town hall (distance by the roads)”, “Aggregate: Year built & Number of storeys”, and “Area of flat”. It can be explained by the fact that Vladivostok is a highly diverse city with an entwinement of historical and modern buildings and RuleFit tries to differentiate them. This is the typical example where linear regression performs poorly and tackling non-linear dependencies is required.
In order to evaluate the effectiveness of RuleFit prediction we can turn to Figure 22 (b) where the comparison of predicted and real values of per square meter price is shown. If we compare Figure 22 (a) and Figure 22 (b) we can see that the cloud of points for Rulefit is more elongated along the diagonal line. It means that the residual variance of RuleFit is less than for OLS regression. Additionally we can note that RuleFit still underestimates high prices but performs better for low prices.
Unfortunately, RuleFit like any other model has its drawbacks. First, as one can see in Table 6, RuleFit sometimes can keep strange rules that are very similar to each other. This is a drawback of the generation of decision trees by the algorithm and it complicates the interpretation of the model. Second, due to the stochastic nature of the tree’s generation process, RuleFit each time will generate new rules. This problem can be solved by setting the state of the random number generator. Additionally the same randomness will lead to a different set of final rules after changing the training data for the model. Thus the RuleFit is a badly reproduced model and because of it, we evaluate the final results of the model on the whole set of data without splitting it on test and train. But in the process of study we did such splitting and made crossvalidation which gives the average to be 0,58. We present the result of the RuleFit model in Table 5.
Average approximation error | ||
0,6 | 13 518,28 | 8,8% |
From Table 5 we can see that the coefficient of determination is equal to 0,6 which is acceptable for our purposes. Values of MAE and average approximation error of RuleFit are lower than for linear regression.
Therefore we can prefer RuleFit for prediction of the per square meter price. Despite its drawbacks mentioned above the RuleFit model fairly well describes the data while to a large extent saving the interpretability.
6 Conclusion
This paper is effort to show the problems that an analyst could encounter in process of modeling the property market and simultaneously to suggest various solutions for overcoming these problems – the methods that can help in working with real market datasets and improve model performance. All stages of building models are described on real data for noncommercial land parcels and flats in Vladivostok city, Primorye region, Russia.
The market data taken from sale offers on the Internet, and the data about the deals between seller and buyer could include mistakes and fictitious values. An analyst should find such cases using various visual and statistical tools. It is necessary to exclude outliers otherwise the predictive model will be incorrect and predictions will be far from the truth. This usually leads to the exclusion of a significant part of data however each step has logical and statistical justifications. If one models the common trend in data then reducing the fluctuations makes the process of modeling simpler. However if one rather tries to take into account detailed characteristics of objects instead of massive evaluation then presented methods should be used carefully.
The step of identification and formation of pricing variables (features) could be a labor-consuming process. But in our view, an analyst shouldn’t restrict yourself only to trivial factors, it is worth trying to create aggregative factors or use methods from related sciences.
The understanding of prerequisites for modeling is crucially important. In this paper, we had a strong requirement for the interpretability of the model. Because of this reason, in the first instance we try to use a linear regression. But even despite the inclusion of spatial features, this algorithm is not always effective. We can see this in the example of the segment “Flat” – some variables are not linearly dependent; assumptions of linear regression are not realized. In such a situation, we have to try alternative modeling practices. Experience and methods from related areas of science and the development of machine learning help to improve the quality of predictions and result in the decrease of errors in the model. In this paper, we apply the hybrid approach by joining together the classical method of building linear regression with geostatistics methods namely the interpolation by kriging. Indeed, there are several premises for doing this. Geostatistics works well with spatial random variables while linear regression allows to highlight the trend in data leaving normally distributed random residuals. The modeling of residuals by kriging allowed to significantly decrease the total error of the model. Another example of the unification of several approaches is “RuleFit” which is a hybrid of well-proven models. The base is the linear regression penalizing for the complexity of the model while additional features are generated by the Random Forest algorithm. Such composition allowed to a considerable extent improve the prediction and the quality of the model, taking into account the non-linear nature of dependencies and also latent factors. On the basis of our work, the implementation of RuleFit looks like a reasonable practice in a situation where several objects (flats) correspond to the unique spatial point (multi-unit residential building) and the geostatistical approach encounters serious difficulties. The successful combination of different methods in one model indicates a good perspective of further unification of well-accepted models in order to improve the efficiency of an interpretable model on complex data.
Appendix A The resulting model for RuleFit
The final model for RuleFit is given by:
where are coefficients assigned by RuleFit; - factor or rule from Table 6 and the naming of factors is given below.
: Area of flat
: Distance to the water objects (straight-line distance)
: Distance to the seacoast (straight-line distance)
: Distance to the business activities center (straight-line distance)
: Distance to the closest railway station (straight-line distance)
: Number of cafes (within a radius of 1 km)
: Distance to the town hall (distance by the roads)
: Aggregate: Year built & Number of storeys
: Wall material: bricky
: Wall material: monolithic
: Storey of flat: the first
: Storey of flat: the last
№ | Factor/Rule | Type | Coefficient |
1 | linear | 6713,23 | |
2 | linear | -438,15 | |
3 | linear | -3,14 | |
4 | linear | 6482,54 | |
5 | linear | -2,83 | |
6 | linear | 7435,15 | |
7 | linear | 2,02 | |
8 | linear | 76,72 | |
9 | linear | -0,24 | |
10 | linear | -2771,67 | |
11 | linear | -387,96 | |
12 | linear | -0,08 | |
13 | rule | -26328,13 | |
14 | rule | 29815,42 | |
15 | rule | -6277,76 | |
16 | rule | 17615,75 | |
17 | rule | 13697,61 | |
18 | rule | -6846,8 | |
19 | rule | 8955,08 | |
20 | rule | -4897,04 | |
21 | rule | -4758,11 | |
22 | rule | -7064,57 | |
23 | rule | -3785,21 | |
24 | rule | 4401,15 | |
25 | rule | 6457,31 | |
26 | rule | -3693,24 | |
27 | rule | 2686,18 | |
28 | rule | 5532,75 | |
29 | rule | 4444,61 | |
30 | rule | -898,74 | |
31 | rule | 718,14 | |
32 | rule | -387,44 | |
33 | rule | 399,47 | |
34 | rule | 16,4 |

### v2

@@
-Modern approaches to building effective interpretable models of the property market using machine learning
+Modern approaches to building interpretable models of the property market using machine learning on the base of mass cadastral valuation
 Abstract
@@
 The main source of this is the huge difference between noisy real market data and ideal data which is very common in all types of tutorials on machine learning.
-This paper covers all stages of modeling: the collection of initial data, identification of outliers, the search and analysis of patterns in data, the formation and final choice of price factors, the building of the model, and the evaluation of its efficiency.
+This paper covers all stages of modeling: the collection of initial data, identification of outliers, the search and analysis of patterns in the data, the formation and final choice of price factors, the building of the model, and the evaluation of its efficiency.
 For each stage, we highlight potential issues and describe sound methods for overcoming emerging difficulties on actual examples.
@@
 Therefore we suggest linear regression with automatic generation and selection of additional rules on the base of decision trees, so called the RuleFit method.
-Thus we show, that despite the strong restriction as the requirement of interpretability which is important in practical aspects, for example, legal matters, it is still possible to build effective models of real property markets.
+Thus we show, that despite such a strong restriction as the requirement of interpretability which is important in practical aspects, for example, legal matters, it is still possible to build effective models of real property markets.
 keywords:
@@
 The problems faced by researchers are the insufficiency of data; the market of interest is undeveloped; explanatory factors are insufficient for prediction of the real estate value and so on.
-There are also problems with the nonlinear influence of some variables on others or the low correlation between explanatory variables and dependent variable.
+There are also problems with the nonlinear influence of some variables on others or the low correlation between explanatory variables and a dependent variable.
 All of the above complicate the prediction process and make it nontrivial.
@@
 Cadastral valuation is a process of estimation of the cadastral value of property to create a clear and valid system of taxation and other relations governed by law.
-In the case of land parcels it could be a calculation of rental charge for land parcel use, the definition of land redemption value or buying from public domain or municipal property; for buildings and building spaces – calculation of taxes for buy/sell, rental and exchange of object, registration of succession rights, definition of exchange value, registration of gift agreement.
+In the case of land parcels it could be a calculation of rental charge for land parcel use, the definition of land redemption value or buying from public domain or municipal property; for buildings and building spaces – calculation of taxes for buy/sell, rent and exchange of an object, registration of succession rights, definition of exchange value, registration of gift agreement.
 The process of determination of cadastral value requires taking into account various factors, related to location, size, and quality of real estate property and also expert assessment of its potential for use in various purposes (Byrda, 2025).
@@
 The precise cadastral valuation promotes the creation of a true picture of real market property values and leads to the improvement of the investment environment and stimulation of the outlays.
-Incorrect cadastral valuation or ineffective using results of cadastral valuation could reduce the investment attractiveness and give rise to a growth rate reduction of local economies and a decline of the social and economic situation.
-Overcharge of territories by unreasonably high values of cadastral value takes a toll on the interests of potential investors and may cause secular stagnation in the property market.
+Incorrect cadastral valuation or ineffective usage of results of cadastral valuation could reduce the investment attractiveness and give rise to a growth rate reduction of local economies and a decline of the social and economic situation.
+Overcharge of territories by unreasonably high cadastral values takes a toll on the interests of potential investors and may cause secular stagnation in the property market.
 Non-optimal instruments of cadastral valuation ignoring the physical-geographic factor of pricing instigates maldistribution of tax burden and as a result, the insufficient fundraising for covering municipalities’ needs.
@@
 The comparative method is typically used in the case when the market of real estate property is highly developed and there is enough information about deals and offers.
-This method includes method of fiducial object and models of regression analysis such as linear regression and its derivatives (Ministry for Economic Development of Russia, 2018).
-However, in the modern world, where technologies improve every day and machine learning actively penetrates all spheres of life, making predictions more and more precise, it becomes clear that we need more widely implement such methods in the mass valuation of real estate to improve the existing algorithms.
+This method includes the method of a fiducial object and models of regression analysis such as linear regression and its derivatives (Ministry for Economic Development of Russia, 2018).
+However, in the modern world, where technologies improve every day and machine learning actively penetrates all spheres of life, making predictions more and more precise, it becomes clear that machine learning algorithms should be more widely implemented in the mass valuation of real estate to improve the existing methods.
 Today there are many scientific works on the application of machine learning to prediction of property value.
@@
 For overcoming this problem authors suggest various spatial methods (Jahanshiri et al., 2011).
-But if models are created to form the base for taxing then transparency becomes the key aspect of the model.
+But if the model is created to form the base for taxing then transparency becomes its key aspect.
 In this context transparency means interpretability.
-The results of the model should be presented in an accessible for ordinary person format – to explain clearly how input data influences on final results.
+The results of the model should be presented in an accessible for an ordinary person format – to explain clearly how input data influences on final results.
 In most cases, the machine learning models are black box – the programmer and users see the input data and the results but the mechanism of calculation is rather implicit.
-Appraiser can more deeply understand these calculation principles and algorithms, and what methods are used for the adjustment of estimates, but explaining it to an ordinary person who only wants to know the answer to just one question “How was the value of my property calculated?”, becomes almost impossible task.
-The interpretability of the model allows to find patterns and estimate how factors influence the price.
+Appraiser can understand these calculation principles and algorithms, and what methods are used for the adjustment of estimates more deeply, but explaining it to an ordinary person who only wants to know the answer to just one question “How was the value of my property calculated?”, becomes almost impossible task.
+The interpretability of the model allows to find patterns and to estimate how factors influence the price.
 It gives an opportunity to explain the constitutes of price to citizens who express disagreement with calculated cadastral values.
@@
 Nowadays there are many reference materials on machine learning, but most of them demonstrate concepts on the base of ideal data.
-These data significantly differ from real data that is subject to errors and fluctuation.
-As a result, when a researcher tries to practice ideas and methods from a book, he is faced with difficulties – either methods don’t work at all or work incorrectly.
+These data significantly differ from real data that is subject to errors and fluctuations.
+As a result, the implementation of the ideas and methods from a book becomes challenging - they either do not work at all or work incorrectly.
 For solving this problem a researcher should undertake careful preparatory work with data.
@@
 For each segment, we analyzed the data and built the model.
-Sections 3 “Methods of identification of the outliers” and 4 “The formation and selection of explanatory variables” present the basic principles and interesting facts of the research performed for both segments.
+In section 2 “Description of the target of research” we briefly introduce the territory for which we build the models and shortly describe the data for modeling.
+More detailed analysis of data is given in the following sections.
+Sections 3 “Identification of outliers” and 4 “Formation and selection of explanatory variables” present the basic principles and interesting facts of the research performed for both segments.
 The section 5 “Modeling” consists of two parts describing each market segment.
 This section presents a detailed description of the methods and principles of modeling.
-2 Description of target of research
+In section 6 “Conclusion” we outline the research and highlight the most important points.
+2 Description of the target of research
 The modeling territorially covers Primorsky Krai – the administrative region of the Russian Federation located in the southeast of the country (Figure 1).
 The region borders China on the west, North Korea on the south-west, and Khabarovsk Krai (Russian Federation) on the north.
-Primorsky Krai is the part of Far Eastern Federal District.
+Primorsky Krai is the part of the Far Eastern Federal District.
 The area of the region is 164 673 km2 which is comparable to areas of such countries as Bulgaria (110 994 km2), Greece (131 957 km2) or Cambodia (181 035 km2).
@@
 In 2012 Vladivostok was the venue for the APEC (Asia-Pacific Economic Cooperation) summit.
-For this event a wide range of significant measures was taken – the development of telecommunication, the construction of new roads, enhancing the cultural and prestigious status of the region (building of Opera and Ballet Theatre, building of multipurpose health center), the reconstruction of energy and heat supply facilities contributing to the reduction of energy costs of enterprises and energy losses in public utilities.
-Vladivostok is one of ten cities in the country where the Federal University was created (its new campus on Russky Island was the venue for APEC).
+For this event a wide range of significant measures was taken – the development of telecommunication, the construction of new roads, the enhancement of the cultural and prestigious status of the region (building of the Opera and Ballet Theater, building of multipurpose health center), the reconstruction of energy and heat supply facilities contributing to the reduction of energy costs of enterprises and energy losses in public utilities.
+Vladivostok is one of ten cities in the country where the Federal University was created (its new campus on the Russky Island was the venue for APEC).
 This university has become a powerful intellectual base for the region.
 The membership of Russia in APEC and Primorsky Krai in particular as the main territory for interaction with other countries of the Asia-Pacific region allows balanced management of the regional economy and effective use of the natural resource potential of the Far East of Russia.
-This in turn increases the attractiveness of the region as a place for residence and creates opportunities for enhancing the business availability and for recruitment of potential investors (Chebotarev, 2012).
+This in turn increases the attractiveness of the region as a place for residence and creates opportunities for enhancement of the business availability and for recruitment of potential investors (Chebotarev, 2012).
 Additionally, Asia-Pacific Economic Cooperation favors regional economic integration, supports a multilateral trading system, unlocks the potential of the digital economy, and helps to bridge the digital divides (Ministry for Economic Development of Russia, ).
@@
 Because of this reason, the results of cadastral valuation must show the real market situation, create a valid taxation system, and be a base for other relations governed by law.
-Precisely because of its favorable territorial location Vladivostok been in a third ten of Russia’s cities by population and separated by thousands of kilometers from the central part of Russia stably occupies a place in the top ten of cities of Russia by real property value.
+Precisely because of its favorable territorial location Vladivostok being in the third ten of Russia’s cities by population and separated by thousands of kilometers from the central part of Russia stably occupies a place in the top ten of cities of Russia by real property value.
 But the modeling of real property value is a difficult task due to several factors: complex terrain, historico-geographical urban development, various transport accessibility of different parts of the city, the existence of insular territories, the mixing of historic and modern buildings, nonuniform density of social, cultural and entertainment centers, coexistence of industrial estate and domestic development and so on.
@@
 The object for predictions is the real estate value, specifically – the Per Square Meter Price of real estate value (PSMP):
+In practice, analysts prefer to model the PSMP but not the total estate price.
+The reason is that the “Area” is fairly significant pricing variable and non-inclusion it to the model will often result in decrease of quality of the model.
+But if one considers the total value of estate price and simultaneously includes the “Area” into a list of independent variables then it will just completely suppress other variables with no contribution to the predictive power of the model.
 Basic data for the real estate value were taken from open sources (like websites with offers for the sale of property) and deals.
@@
 It’s worth noting that in the process of analyzing the data on the outliers, one should take into account the type of objects.
-A land parcel is basically a one-dimensional object.
-Usually, the price of such an object depends on its distance relative to other objects, for example center of the town.
-On the other hand, the value of a flat depends not only on its location but also on the characteristics of the multi-unit residential building where the flat is located.
-In the second case, the process of identification of outliers can be more difficult.
-In the case when we analyze the one-dimensional objects (land parcels), we can plot their per square meter price on a map and estimate their validity from general considerations.
+A land parcel loosely could be considered as an one-dimensional object.
+We mean that basically the geographic coordinates of a land parcel define its price.
+Coordinates decode both the spatial location of the object and its individual characteristics such as the slope, the availability of electricity, and so on.
+There is one to one mapping between coordinates and the land parcel.
+On the other hand, the typical multi-unit residential building contains several tens or even hundreds of flats which means that one spatial point corresponds to multiple objects.
+Flats in the same building could differ by the storey or the area.
+Even the adjacent buildings having almost the same coordinates could differ drastically by the year of built, the wall material, etc.
+This is fairly common situation for flats and rather untypical for land parcels where the neighboring objects tend to have similar characteristics.
+As the result, the process of identification of outliers for the former could be more difficult.
+For the latter case, we can plot the per square meter price of land parcels on a map and estimate its validity from general considerations.
 Figure 2 shows that the points located close to each other can have significantly different per square meter prices.
@@
 All these factors make land parcel exploitation more difficult.
-However, the determination of the cadastral value of the object is the mass process reflecting the value of a typical object without delving too much into its specific characteristics.
-Because of this, reason we can expect a similar per square meter price among land parcels in some local area that changes with distances to other objects included in price factors (center of a town, water objects, social objects, etc.).
+However, more likely, a such difference in prices between neighboring land parcels means that the one or both prices are anomalous and should be considered as outliers.
+The price of objects like land parcels tends to change rather continuously with distance between them without sharp variations.
+Because of this reason, we can expect a similar per square meter price among land parcels in some local area that changes with distances to other objects included in price factors (center of a town, water objects, social objects, etc.).
 For the exclusion of outliers, we can start with the clusterization of points on the base of their spatial relationship.
-The next thing we can calculate the median or average value for each cluster and find the outliers using such methods as box-plot or z-score or so on.
+Next we can calculate the median or average value for each cluster and find the outliers using such methods as box-plot or z-score and so on.
 Also, such clustering helps to take into account the inhomogeneity of administrative territory.
 For example, the territory could include the city and the suburbs where a per square meter price is sufficiently different (Figure 3).
-On the contrary, if we consider the whole territory without clusterization we can lose part of the data because firstly the per square meter price of the city is much higher than the per square meter price of the suburbs and secondly the number of data points in the city are much less than the corresponding number in suburbs.
+On the contrary, if we consider the whole territory without clusterization we can lose part of the data because firstly the per square meter price of the city is much higher than the per square meter price of the suburbs and secondly the number of data points in the city is much less than the corresponding number in suburbs.
 In this case, the land parcels located in the city would be outliers and by excluding them we will have major losses of important data and our predicted value would be sufficiently distorted.
-Also, we can find the outliers using the probability density function of dependent variable. For example, Figure 4 shows that values between 50000 rub and 100000 rub for the deals are doubtful.
-We can confirm this assumption by taking buildings and analyzing per square meter prices in each building.
+Also, we can find the outliers using the probability density function of the dependent variable. For example, Figure 4 shows that values between 50000 rub and 100000 rub for the deals are doubtful.
+We can check this assumption by taking buildings and analyzing per square meter prices in each building.
 The example of one building is shown in Table 1 where suspicious value is highlighted.
@@
 Type | Storey of flat | Total price, rub | Area of flat, m2 | PSMP, rub |
-offer | 5 | 3 850 000 | 17,1 | 225 146 |
-offer | 2 | 4 500 000 | 23,1 | 194 805 |
-offer | 8 | 6 150 000 | 34,5 | 178 261 |
-offer | 9 | 4 350 000 | 17,0 | 255 882 |
-offer | 3 | 4 200 000 | 17,0 | 247 059 |
-offer | 4 | 3 800 000 | 16,7 | 227 545 |
-offer | 5 | 3 800 000 | 17,0 | 223 529 |
-deal | 5 | 4 300 000 | 16,7 | 257 485 |
-deal | 2 | 4 100 000 | 23,0 | 178 261 |
-deal | 4 | 4 050 000 | 16,9 | 239 645 |
-deal | 4 | 900 000 | 16,8 | 53 571 |
-deal | 4 | 4 900 000 | 23,3 | 210 300 |
-deal | 4 | 3 800 000 | 16,7 | 227 545 |
+offer | 5 | 3 850 000 | 17.1 | 225 146 |
+offer | 2 | 4 500 000 | 23.1 | 194 805 |
+offer | 8 | 6 150 000 | 34.5 | 178 261 |
+offer | 9 | 4 350 000 | 17.0 | 255 882 |
+offer | 3 | 4 200 000 | 17.0 | 247 059 |
+offer | 4 | 3 800 000 | 16.7 | 227 545 |
+offer | 5 | 3 800 000 | 17.0 | 223 529 |
+deal | 5 | 4 300 000 | 16.7 | 257 485 |
+deal | 2 | 4 100 000 | 23.0 | 178 261 |
+deal | 4 | 4 050 000 | 16.9 | 239 645 |
+deal | 4 | 900 000 | 16.8 | 53 571 |
+deal | 4 | 4 900 000 | 23.3 | 210 300 |
+deal | 4 | 3 800 000 | 16.7 | 227 545 |
 More serious methods such as machine learning also can be used for the identification of outliers.
-Particularly it would help for many-dimensional objects e.g.
-flats.
+Particularly it would help for “multi-dimensional” objects like flats.
 Firstly we should perform territorial clustering of points that results in the division of the populated place into regions (clusters).
 Next, we create the new clusters in each region on the base of the characteristics of the buildings e.g.
-“Wall material”, “Number of stories” and “Year built”.
+“Wall material”, “Number of storeys” and “Year built”.
 This procedure allows to highlight typical buildings.
@@
 Such algorithms take as input the buildings in the space of the above-mentioned characteristics and return the class markers for each building.
-In this paper, we use the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996).
+In this paper, we use the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm (Ester et al., 1996) with realization on Python (the library “Scikit-learn” (Pedregosa et al., 2011)).
 This algorithm is based on the idea of closeness between points in the space of features.
@@
 A model that contains a smaller number of outliers is preferable.
-In our case, we can transform the exponential dependence to linear by taking the natural logarithm of data.
+In our case, we can transform the exponential dependence to linear by taking the natural logarithm of the data.
 So for each cluster found by DBSCAN, we apply RANSAC regression to find the outliers and to highlight the trend.
-The dependent variable is the natural logarithm of the per square meter price and the explanatory variable is the area of flat.
+The dependent variable is the natural logarithm of the per square meter price and the explanatory variable is the area of a flat.
 This procedure allows to find the trend for data with high fluctuations and noise and hinges on the reasonable assumption that flats in buildings with similar characteristics located within some compact territory should have similar per square meter prices.
@@
 The example of the results of RANSAC regression is shown in Figure 5, the points marked as outliers are not used for forecasting.
+We used median absolute deviation (MAD) as the threshold.
+The “SciKit-learn” library uses it by default.
 4 Formation and selection of explanatory variables
-According to methodological rules of mass valuation of real estate, we can use for explanatory variables such information as general characteristics of an object (object’s purpose, area, object’s name, wall material, year build, extension of the land parcel, kind of soil and so on); information about a location of an object (a distance to the roads, to the water objects, to a recreation area, a city center and so on); information about engineering infrastructure (a class of linear object, a distance to a power line, a rated power of a linear object and so on); other matter (a price level of a commodity bundle, commodity circulation per 1 person, a distance to extractive industry territory) (Ministry for Economic Development of Russia, 2018).
+According to methodological rules of mass valuation of real estate, we can use as explanatory variables such information as general characteristics of an object (object’s purpose, area, object’s name, wall material, year built, extension of the land parcel, kind of soil and so on); information about a location of an object (a distance to the roads, to the water objects, to a recreation area, a city center and so on); information about engineering infrastructure (a class of linear object, a distance to a power line, a rated power of a linear object and so on); other matter (a price level of a commodity bundle, commodity circulation per 1 person, a distance to extractive industry territory) (Ministry for Economic Development of Russia, 2018).
 General information about the object is contained in the Unified State Register of Real Estate and in accordance with the law is given by the ministry to the certified organization which makes evaluation.
@@
 Geographical features can be extracted with the help of mapping software – NextGIS and ArcGIS.
-Also, we can get such information using programming tools, for example, Python and open non-profit mapping source – OpenStreetMap.
-Using the library “osmnx” from Python we can get all kinds of socially significant facilities – nursery schools, schools, hospitals, shops, malls, and so on.
+Also, we can get such information using programming tools, for example, Python and open non-profit mapping source – OpenStreetMap (OpenStreetMap contributors, 2017).
+Using the library “OSMnx” (Boeing, 2025) from Python we can get all kinds of socially significant facilities – nursery schools, schools, hospitals, shops, malls, and so on.
 Then we can calculate the straight-line distances to these objects or, for example, count the number of such objects in some radius.
 We can use this data as explanatory variables for a predictive model.
-Also libraries “osmnx” and “networkx” allow loading road network graphs.
+Also libraries “OSMnx” and “NetworkX” (Hagberg et al., 2008) allow the loading of road network graphs.
 So we have the opportunity to calculate the distance by the roads.
 This calculation improves the accuracy of the variables.
-It can be important when we have a territory with irregular coastlines (Figure 6) or deal with islands.
+It can be important when we have a territory with irregular coastlines (Figure 6), or mountain ranges, or deal with islands.
 In the last case, we can use adjusting factors reflecting the accessibility of the island.
 Also, we can aggregate some variables into one (as explained below in the discussion of the correlation matrix) or use conceptions from the related fields, for example, graph theory.
-Thus we created the factor “Development of the road network” that helped to estimate the location of object in relation to transport accessibility.
+Thus we created the factor “Development of the road network” that helped to estimate the location of an object in relation to transport accessibility.
 If the object has many crossroads and roads nearby then this object is more attractive and therefore more expensive.
@@
 In this paper, we use harmonic centrality (Marchiori and Latora, 2000) because it is simple for calculations and understanding.
-The harmonic centrality of a node is the sum of inverse lengths of the shortest routes for each node to a particular node.
+The harmonic centrality of a node is the sum of inverse lengths of the shortest routes from each node to this node.
 It has a perfect intuitive interpretation – smaller distances give a higher contribution to centrality.
 Mathematically the harmonic centrality can be written as:
-where - the node for which the centrality is calculated; - the other nodes of the graph; - the shortest route between and .
+where is the node for which the centrality is calculated; is the other nodes of the graph; is the length of the shortest route between and .
 The shortest route between two points is the route for which the sum of weights of constituting edges is minimal.
 The roads can have bi-directional and unidirectional traffic.
-Because the graph of the road network is an oriented graph, it means that every edge of the graph has a direction.
+Because the graph of the road network is an oriented graph, it means that every edge of the graph has a direction so the inversion of end points of the route potentially could change its length.
 We should take this fact into account when calculating the centrality.
@@
 first step: find the closest node of the road network to the object;
-second step: consider this result as an estimation of the centrality of the object. -
+second step: consider its centrality as an estimation of the centrality of the object. -
 2.
@@
 In this method, we get estimation as a linear combination of basis functions (Dem’yanov and Savel’eva, 2010; Buhmann, 2000; Regis and Shoemaker, 2006; Zhang et al., 2024).
-where - estimation at the point ; - distance from point to point ; - basis function that depends on the distance; - weight coefficients.
+where is the estimation of centrality at the point ; is the distance from point to point ; is the basis function that depends on the distance; are weight coefficients.
 Coefficient defines algebraic sign of an element and its contribution.
@@
 The weight coefficients are derived from the condition that at known points the results of interpolation must coincide with known values, i.e.
-the interpolation model in all points must estimate values equal to .
+the interpolation model in all points must provide values equal to .
 Thus we need to solve the system of linear equations with unknown variables.
-To do this we use the function RBFInterpolator from Python library ”SciPy”.
+To do this we use the function RBFInterpolator from Python library “SciPy” (Virtanen et al., 2020; SciPy team, ).
 The results of interpolation for Vladivostok are shown in Figure 7 (b).
 Here for the centrality of the objects we use the term “Development of the road network” underlining our mathematical interpretation of this concept.
+(a) | (b) |
 If we want to build a model for an inhomogeneous area (for example if territory contains a few towns or villages) then we can try to introduce new variables, using aggregation of existing variables.
-For example, we created the factor “Population of a residential place divided on the distance to Vladivostok by the roads”.
+For example, from Figure 8 (a) one can see that independent variables “: Population of a human settlement” and “: Distance to Vladivostok (by roads)” have quite significant correlation with the dependent variable “: Per square meter price” around 0.5.
+But we can achieve even higher correlation with by creating the new factor “Population of a residential place divided by the distance to Vladivostok by roads”.
 On the one hand, the increasing of distance between the residential place and Vladivostok’s administration negatively affects the per square meter price; on the other hand, the population of the residential place positively affects the per square meter price.
-These considerations led us to create the factor that aggregates two variables – the population of a residential place and the distance to Vladivostok by the roads.
-Econometrically this factor should have a positive correlation with the per square meter price – with an increase in the population of residential place and a decrease inf the distance to Vladivostok the per square meter price increases.
-Figure 8(a) shows the relation between this factor and the average value of per square meter price over a building.
-We can notice that data contains positive a linear correlation however we also can notice that there are quite a lot of points when the aggregate factor has low values.
-Because of this, we tried to raise this factor to different powers.
-Figure 8(b) shows the relation between this factor raised to 1/3 power and the average value of per square meter price over a building.
-Such transformation gives a better value of correlation.
-(a) | (b) |
+These considerations led us to create the factor that aggregates two variables – the population of a residential place and the distance to Vladivostok by roads.
+Econometrically this factor should have a positive correlation with the per square meter price – with an increase in the population of residential place and a decrease in the distance to Vladivostok the per square meter price increases.
+Figure 8 (b) shows the relation between this factor and the value of per square meter price of flat.
 The next important step in the way of creating the model is the estimation of correlations of all factors with the dependent variable (per square meter price) and the selection of the most important of them.
 For this step, we created the matrix of correlation for quantitative factors and we analyzed probability density functions of per square meter price for categorical factors.
-When we study the matrices of correlation (Figure 9, Figure 10) we are interested not only in the relationship between explanatory variables and dependent variable but also in the relationships between explanatory variables with each other.
+When we study the matrices of correlation (Figure 9, Figure 10) we are interested not only in the relationship between the explanatory variables and the dependent variable but also in the relationships between explanatory variables.
 The first model that we tried was the linear regression, because of its interpretability.
 The linear model has a set of assumptions, one of them is the absence of multicollinearity in data.
-Multicollinearity is a situation when two or more variables have a significant correlation with one another ().
+Multicollinearity is a situation when two or more variables have a significant correlation with one another ( (Kozak, 2009; Gujarati and Porter, 2009)).
 In this case, the inclusion of them in the model becomes unpractical.
@@
 Previously we noted that before including categorical factors into a model we plot probability density functions of a per square meter price.
-A similar method of analysis would be drawing of a box-plot diagram.
+A similar method of analysis would be drawing of a box-plot diagram (Tukey, 2000).
 Traditionally the categorical factors are included in a model as binary features and one should pay attention not only to average values of per square meter price for each group but also to the representativity of the group in the data.
 For example, Figure 11 shows the probability density function for the factor “Wall material” and we see that flats in timber buildings have a per square meter price far less than the other materials.
-For this reason, we can’t include this binary feature in a model due to its nonrepresentativity.
-A possible solution to this problem can be the introduction of a correction factor reducing the per square meter price of such objects after getting of estimation of a model.
+For this reason, we can’t include this binary feature in a model due to its non-representativity.
+A possible solution to this problem can be the introduction of a correction factor reducing the per square meter price of such objects after getting of the estimation of a model.
 After completing all preparative steps – removing multicollinearity; dropping the factors without the economic substance; and dropping the non-significant features, we can move to the question of selection of price-forming factors that we may include in the model.
-There are different ways to solve this problem (Ajvazyan, 2010) – we can try all possible combinations of factors and choose the best model taking into account some requirements (minimizing errors or maximizing ); we can choose a model that meets the initial conditions for its using (for example for ordinary least squares such conditions are unbiasedness, consistency and effectiveness of estimations); also we can use algorithms for automatic feature selection – Forward selection or Backward elimination (Narisetty, 2020).
-For segment “Land parcels” where the correlation of features with dependent variable has sufficiently high values (Figure 9) we applied Forward selection of features and also we took into account the value of and realization of the basic conditions for ordinary least squares method (Ajvazyan, 2010; Kutner, 2005) (Durbin–Watson statistic, Jarque-Bera test, F-test).
-Number of factors | D–W stat | J-B (prob) | F-stat (prob) | |
-1 | 0,636865 | 1,238017 | 0,000005 | 1,173483e-207 |
-2 | 0,712070 | 1,371189 | 0,072956 | 3,926066e-253 |
-3 | 0,749263 | 1,564183 | 0,175133 | 1,023501e-279 |
-4 | 0,752834 | 1,564116 | 0,073893 | 2,622868e-281 |
-5 | 0,763181 | 1,609770 | 0,002961 | 1,090438e-288 |
-6 | 0,773200 | 1,566779 | 0,004079 | 3,415025e-296 |
-From Table 2 one can see that in this particluar case, 4 factors are the optimal choice for modeling linear regression taking into account the conditions for using ordinary least squares.
+There are different ways to solve this problem (Ajvazyan, 2010) – we can try all possible combinations of factors and choose the best model taking into account some requirements (minimizing errors or maximizing ); we can choose a model that meets the initial conditions for its usage (for example for ordinary least squares such conditions are unbiasedness, consistency and effectiveness of estimations (Gujarati and Porter, 2009)); also we can use algorithms for automatic feature selection – Forward selection or Backward elimination (Narisetty, 2020).
+For the segment “Land parcels” where the correlation of features with dependent variable has sufficiently high values (Figure 9) we applied Forward selection of features and also we took into account the value of (Gujarati and Porter, 2009) and realization of the basic conditions for ordinary least squares method (Ajvazyan, 2010; Kutner, 2005): residuals should not have the significant autocorrelation which corresponds to the Durbin–Watson statistic to be in range from 1.5 to 2.5 where the value 2.0 means that residuals are completely uncorrelated (Field, 2012); residuals should be normally distributed which means they should pass the Jarque-Bera test with -value(JB) 0.05 (Gujarati and Porter, 2009) that corresponds to do not rejecting the null hypothesis of normality; the equation of regression should be significant with -value(-statistic) 0.01 (Gujarati and Porter, 2009) meaning we should reject the null hypothesis that all coefficients of the model except the intercept are zero.
+Number of factors | Durbin-Watson | -value(Jarque-Bera) | -value(-statistic) | |
+1 | 0.634 | 1.238 | 0.000005 | 1.173483e-207 |
+2 | 0.712 | 1.371 | 0.073 | 3.926066e-253 |
+3 | 0.749 | 1.564 | 0.175 | 1.023501e-279 |
+4 | 0.753 | 1.564 | 0.074 | 2.622868e-281 |
+5 | 0.763 | 1.610 | 0.003 | 1.090438e-288 |
+6 | 0.773 | 1.567 | 0.004 | 3.415025e-296 |
+From Table 2 one can see that in this particular case, 4 factors are the optimal choice for modeling linear regression taking into account the conditions for using ordinary least squares.
+The another option is to select 3 factors - it also satisfies the statistical tests, has the similar .
+But all other things being equal we think it is better to keep more factors.
 5 Modeling
@@
 All these territories have different property markets, price levels, and quality of life, that depend on economic and territorial characteristics.
-The methodological rules (Ministry for Economic Development of Russia, 2018) recommend splitting the territory on the base of administrative boundaries of regions taking into account the population.
+The methodological rules (Ministry for Economic Development of Russia, 2018) recommend splitting the territory on the base of administrative boundaries of municipalities taking into account the population.
 This recommendation can be applied to objects that are located on homogeneous territory, for example for flats that are located in one city.
@@
 land parcels.
-They occupy a large heterogeneous area – the city and the suburbs and often form allotment societies (also called gardeners’ non-commercial partnerships).
+They occupy a large heterogeneous area – the city and the suburbs, and often form allotment societies (also called gardeners’ non-commercial partnerships).
 Gardeners’ non-commercial partnerships are a collective form of gardening.
@@
 Similar organizations exist in many countries, for example, Germany, Sweden, Poland, and the USA.
-In every country, there are own procedures for the registration and obligation to pay a tax.
-In Russia, such partnerships must pay taxes, in particular land tax (Seregina, 2019) – the tax paid by physical persons or organizations that have property right of land plot or right of land plot perpetual usage or right of lifetime inheritable possession.
-Tax amount is defined from cadastral value and taxation rate.
+In every country, there are its own procedures for registration and obligation to pay taxes.
+In Russia, such partnerships must pay taxes, in particular the land tax (Seregina, 2019) – the tax paid by individuals or legal entities that have the property right of land plot or the right of land plot perpetual usage or the right of lifetime inheritable possession.
+Tax amount is defined from the cadastral value and taxation rate.
 Such an allotment society can include different administrative regions.
@@
 Most of the algorithms are based on spatial relations between objects.
-Because of this reason, for the segment “Land parcels” we used a k-means clustering algorithm (MacQuenn, 1967), that takes into account the spatial positions of objects.
+Because of this reason, for the segment “Land parcels” we used a k-means clustering algorithm (MacQuenn, 1967), that takes into account spatial positions of objects.
 This method is simple for understanding and for realization.
@@
 By using this algorithm we are more confident that neighboring points belong to one cluster.
-All further calculations of modeling were done for the capital region – Vladivostok and its suburbs.
+We used Scikit-learn realization of this algorithm.
+All further calculations were done for the capital region – Vladivostok and its suburbs.
 segment “Land parcels”
+(a) | (b) |
 First of all, we plotted the probability density function of the dependent variable (per square meter price) and noticed that it is not normal (Figure 12 (a)).
-Although statistic and machine learning theories don’t have the strict rule requiring the normality of the probability density function of the dependent variable, usually analysts prefer to work with data distributed approximately normally.
+Although there is no requirement on the normality of the dependent variable, analysts tend to prefer working with the data distributed approximately normally.
 Because of this reason, we used a natural logarithm of the per square meter price (Figure 12 (b)).
 Also such transformation of data guarantees that prediction of modeling will have only positive values.
-(a) | (b) |
+OLS Regression Results | ||||
+Dep. variable: | ||||
+: | 0.761 | |||
+: | 0.760 | |||
+-statistic: | 740.8 | |||
+-value (-statistic): | 1.27-287 | |||
+coef | std err | t | ||
+7.3210 | 0.012 | 627.655 | ||
+0.1070 | 0.014 | 7.585 | ||
+-0.1642 | 0.013 | -12.254 | ||
+-0.3719 | 0.016 | -22.918 | ||
+-0.1153 | 0.009 | -12.876 | ||
+is Natural logarithm of per square meter price | ||||
+is Distance to the water objects (straight-line distance) | ||||
+is Distance to the seacoast (straight-line distance) | ||||
+is Distance to the town hall (straight-line distance) | ||||
+is Aggregating factor: , , (PCA) | ||||
+is Distance to the closest hospital (straight-line distance) | ||||
+is Distance to the closest nursery school (straight-line distance) | ||||
+is Distance to the closest school (straight-line distance) | ||||
+Durbin-Watson: | 1.572 | |||
+Jarque-Bera (JB): | 3.713 | |||
+-value (JB): | 0.156 | |||
+Cond. No.: | 2.87 |
 We start the modeling by trying the simple and interpretable model – the linear regression.
 As we mentioned above for the selection of features we apply Forward selection using a correlation matrix (Figure 9).
-The data was split into two parts – a training set (70% of data) and a testing set (30% of data).
-Figure 13 shows the result of the training of the model.
-Using Figure 13 we can check the model on basic requirements for linear regression.
-We see that the coefficients of variables are significant (); in general, the equation of regression is significant (); the coefficients of variables have the right sign according to economic substance; model residuals have normal distribution () and don’t have autocorrelation (); adjusted coefficient of determination is significant () and it is not differ sufficiently from .
-Also factors included in the model don’t have multicollinearity because the value of the variance inflation factor (VIF) is less than 10 for each factor (Shrestha, 2020) (Table 3).
-Factors | |
-Distance to the water objects (straight-line distance) | 1,463 |
-Distance to the seacoast (straight-line distance) | 1,320 |
-Distance to the town hall (straight-line distance) | 1,936 |
-Aggregate: Distance to the hospitals (straight-line distance) & Distance to the closest nursery school (straight-line distance) & Distance to the closest school (straight-line distance) | 1,424 |
-Figure 14 shows the distribution of model residuals.
+The data was split into two parts – a training set (70% of the data) and a testing set (30% of the data).
+Table 3 shows the result of the training of the model (for building linear regression we used the “statsmodels” (Seabold and Perktold, 2010) library).
+Factors | VIF |
+Distance to the water objects (straight-line distance) | 1.463 |
+Distance to the seacoast (straight-line distance) | 1.320 |
+Distance to the town hall (straight-line distance) | 1.936 |
+Aggregate: Distance to the hospitals (straight-line distance) & Distance to the closest nursery school (straight-line distance) & Distance to the closest school (straight-line distance) | 1.424 |
+Using this table we can check the model on basic requirements for linear regression.
+We see that the coefficients of variables are significant (-value 0.01); in general, the equation of regression is significant (-value(-statistic) 0.01); the coefficients of variables have the right sign according to economic substance; model residuals are normally distributed (-value(JB) 0.05) and do not have the strong autocorrelation (1.5 Durbin-Watson 2.5); adjusted coefficient of determination is significant () and it does not differ sufficiently from .
+Also factors included in the model do not have multicollinearity because the value of the variance inflation factor (VIF) is less than 10 for each factor (Shrestha, 2020) (Table 4).
+Figure 13 shows the distribution of the model residuals.
 This distribution is approximately normal which satisfies the assumptions of linear regression.
 Also, we tested the normality of residuals using the library “SciPy”.
-This test compares calculated with the significance threshold (for this paper we chose ).
-If calculated is more than we can assume that the distribution of model residuals is normal (in our case ).
-Also, we created the scatter plot values of residuals from their numerical order.
+This test compares calculated -value with the significance threshold (for this paper we chose ).
+If calculated -value is more than we can assume that the distribution of model residuals is normal (in our case -value = 0.068) (D’Agostino, 1971; D’Agostino and Pearson, 1973).
+Also, we created the scatter plot of values of residuals from their numerical order.
 From Figure 14 we can see that the data does not contain any dependencies.
@@
 On the base of the described results, we can state that our model is effective and interpretable.
-As we’ve already said the key factor for the identification of the price of a land parcel is its location.
+As we have already said the key factor for the identification of the price of a land parcel is its location.
 We see that linear regression does relatively well but the price of an object can include latent (local-territorial) characteristics.
@@
 There are two basic conditions for using these methods – data should have spatial continuity and stationarity (Kovalevskiy, 2012).
-The first condition means that two points located close together should have rather similar values than points separated more.
+The first condition means that shorter distances between the points should in general lead to closer values of the variable in these points.
 The second condition means that in some locality the studied phenomenon must be spatially homogeneous.
 We can check the condition of stationarity using the probability density function of a variable, in our case we analyze the distribution of residuals of our linear model.
-If the histogram is similar to the Gaussian distribution then in all likelihood there is stationarity (statistical homogeneity).
+If the histogram is similar to the Gaussian distribution then we can assume the stationarity (statistical homogeneity).
 Strictly speaking, this criteria is not a sufficient condition for confirmation of stationarity but it is strong enough and sufficient condition for practical work (Kovalevskiy, 2012).
-Figure 14 and the result of the Jarque–Bera test confirm the stationarity in linear model residuals.
-It’s worth noting that before checking data on stationarity we need to examine whether data contains a trend.
-In our case, we removed the trend from data using linear regression and then we modeled its residuals with geostatistical methods.
-The trend in data can distort the variogram – an instrument for estimation of spatial correlations in data (Dem’yanov and Savel’eva, 2010; Isaaks, 1989).
+Figure 13 and the result of the Jarque–Bera test confirm the stationarity in the linear model residuals.
+It is worth noting that before checking the data on stationarity we need to examine whether the data contains a trend.
+In our case, we removed the trend from the data using linear regression and then we modeled its residuals with geostatistical methods.
+The trend in the data can distort the variogram – an instrument for estimation of spatial correlations in data (Dem’yanov and Savel’eva, 2010; Isaaks, 1989).
 The variogram model is a function describing a dependence between the studied spatial variable and the distance.
@@
 From Figure 16 we see that with increasing distance between points, the correlation decreases.
-Therefore this criteria is satisfied.
-Thus the final value of the per square meter price is the sum of the trend that we got using the linear regression and the prediction of linear model residuals that we got using kriging.
+Therefore this condition is satisfied.
+Thus the final value of the per square meter price is the sum of the trend that we got using the linear regression and the prediction of the linear model residuals that we got using kriging.
 The last step of modeling is getting the resulting predicted estimations – we need to take the exponent of the predicted logarithmic value.
@@
 segment “Flats”
-From the correlation matrix (Figure 10) we can see that correlations between the dependent variable and pricing variables for the segment “Flat” are significantly smaller than for the segment “Land parcels”.
-There is no explicit dependence on the object’s location.
-Nevertheless, we see that socially important factors such as “Number of shopping malls (within a radius of 1 km)” and “Number of cafes (within a radius of 1 km)” have higher correlations with the per square meter price than other factors, but still insufficient.
+OLS Regression Results | ||||
+Dep. variable: | ||||
+: | 0.319 | |||
+: | 0.318 | |||
+-statistic: | 227.8 | |||
+-value (-statistic): | 1.22-317 | |||
+coef | std err | t | ||
+1.538+05 | 378.085 | 406.867 | ||
+-8730.2980 | 384.482 | -22.707 | ||
+3472.7788 | 552.203 | 6.289 | ||
+-1075.3009 | 465.506 | -2.310 | ||
+-3200.5609 | 445.592 | -7.183 | ||
+3374.1161 | 495.161 | 6.814 | ||
+4874.6153 | 439.653 | 11.087 | ||
+-4124.0290 | 455.890 | -9.046 | ||
+-5892.8762 | 384.745 | -15.316 | ||
+is Per square meter price | ||||
+is Area of flat | ||||
+is Distance to the water objects (straight-line distance) | ||||
+is Distance to the seacoast (straight-line distance) | ||||
+is Distance to the closest railway station (straight-line distance) | ||||
+is Distance to the area of negatively influencing factors (straight-line distance) | ||||
+is Number of cafes (within a radius of 1 km) | ||||
+is Distance to the town hall (distance by the roads) | ||||
+is Wall material: large-panel | ||||
+Durbin-Watson: | 1.952 | |||
+Jarque-Bera (JB): | 791.269 | |||
+-value (JB): | 1.51-172 | |||
+Cond. No.: | 2.96 |
+From the correlation matrix (Figure 10) we can see that correlations between the dependent variable and the pricing variables for the segment “Flats” are significantly smaller than for the segment “Land parcels”.
+There is no explicit dependence on the object’ location.
+Nevertheless, we see that socially important factors such as “Number of shopping malls (within a radius of 1 km)” and “Number of cafes (within a radius of 1 km)” have higher correlations with the per square meter price than other factors, although still insufficient.
 However, we start with an effort to create a simple and interpretable model – linear regression.
-This step is necessary according to methodological rules of mass valuation of real estate (Ministry for Economic Development of Russia, 2018).
-Before modeling we did data preprocessing – we selected the factors (features) included in the model and made data standardization (Konayeva et al., 2024).
+This step is necessary according to the methodological rules of mass valuation of real estate (Ministry for Economic Development of Russia, 2018).
+Before modeling we prepared the data by selecting the the factors (features) included in the model and performing the data standardization (Konayeva et al., 2024).
 The standardization is useful in the case of considerably different values of factors in order to reduce them to one scale.
-As we mentioned earlier the selection of the variables can be done in various ways according to theoretic assumptions and/or assumptions of area under the study.
-In this paper we used a machine algorithm of selection – SelectKBest (library “sklearn”).
-This method estimates the dependence of factors with a dependent variable using statistic tests and selects the best factors according to preferred criteria.
-We used the F-test as a statistical test for estimating the linear dependence between factors.
+As we mentioned earlier the selection of the variables can be done in various ways according to theoretical assumptions and/or specific aspects of studied area.
+In this paper we used a machine algorithm of selection – SelectKBest (library “SciKit-learn”).
+This method estimates the influence of the factors on the dependent variable using statistic tests and selects the best factors according to preferred criteria.
+We used the -test as a statistical test to estimate the linear dependence between the factors and the per square meter price.
 We selected 8 the most significant pricing variables using SelectKBest – “Area of flat”, “Distance to the water objects (straight-line distance)”, “Distance to the seacoast (straight-line distance)”, “Distance to the closest railway station (straight-line distance)”, “Distance to the area of negatively influencing factors (straight-line distance)”, “Number of cafes (within a radius of 1 km)”, “Distance to the town hall (distance by the roads)”, “Wall material: large-panel”.
 We created the linear regression model using these factors and taking as the dependent variable the per square meter price without any more transformations.
-The results of the model are shown in Figure 19.
-From Figure 19 we can see that the coefficients of variables are significant (); in general the regression is significant (); the coefficients of variables have right sign according to economic substance; model residuals don’t have autocorrelation (); the value of Condition Number is 2,96, it means that factors don’t contain multicollinearity.
-Also, important conditions for using a linear regression are expected value of residuals must be equal to zero (in our case we got the value 3,61e-10) and the dispersion of residuals must be constant (Figure 20) without any trend.
-But the residuals of our model are not normally distributed () (Figure 21), which means that estimates got by the ordinary least squares are not optimal – they don’t have minimal dispersion in the class of linear unbiased estimators.
+The results of the model are shown in Table 5.
+From Table 5 we can see that the coefficients of variables are significant (-value 0.05); in general the regression is significant (-value(-statistic) 0.01); the coefficients of variables have right sign according to economic substance; model residuals do not have autocorrelation (1.5 Durbin-Watson 2.5); the value of Condition Number (Gujarati and Porter, 2009) is 2.96, it means that factors do not contain multicollinearity.
+Also, important conditions for using a linear regression are the expected value of residuals must be equal to zero (in our case we got the value 3.61e-10) and the dispersion of residuals must be constant (Figure 18) without any trend.
+But the residuals of our model are not normally distributed (-value(JB) 0.05) (Figure 19), which means that estimates got by the ordinary least squares are not optimal – they do not have minimal dispersion in the class of linear unbiased estimators.
 Also in this model we have relatively low values of (), which means that the explanatory power of the model is poor.
 Figure 20 (a) shows the comparison of predicted and real values of per square meter price.
-In an ideal case, the cluster of points is expected to be an ellipse stretched out along line y=x, which means that our predictive values should approach real values with the least dispersion.
+In an ideal case, the cluster of points is expected to be an ellipse stretched out along line y=x, which means that our predicted values should approach real values with the least dispersion.
 But from Figure 20 (a) we can see that in most cases our predictions overestimate the low values and underestimate the high values.
 (a) | (b) |
-We summarized the results of the linear regression model in Table 4. We can use it for comparison with other models.
-train dataset | test dataset | train dataset | test dataset | Average approximation error (train dataset) | Average approximation error (test dataset) |
-0,319 | 0,339 | 17 890,32 | 17 947,46 | 11,7% | 11,9% |
+We summarized the results of the linear regression model in Table 6. We can use it for comparison with other models on the base of the following metrics: , mean absolute error ()
+and average approximation error ()
+train dataset | test dataset | train dataset | test dataset | train dataset | test dataset |
+0.319 | 0.339 | 17 890.32 | 17 947.46 | 11.7% | 11.9% |
 Sometimes the data can include non-linear dependencies that make a linear regression ineffective.
-In such a situation, we can try other algorithms that work better with non-linear dependencies in data.
+In such a situation, we can try other algorithms that work better with non-linear dependencies in the data.
 The most popular method is the “Decision tree” (Díaz-Ramírez et al., 2023).
@@
 This is an advantage of the algorithm.
-But also method has significant shortage – it is prone to overfitting.
+But this method also has a significant shortage – it is prone to overfitting.
 It means that generated rules will fit only the training dataset, this tree will have bad results on other datasets.
@@
 The advantage of it is the high accuracy of prediction and also this algorithm tends to work well with hyperparameters by default.
-Compared to the “Decision tree” the “Random Forest” is much less prone to overfitting.
+Compared to the “Decision tree”, the “Random Forest” is much less prone to overfitting.
 The shortage of this method is the complicated interpretation of modeling results due to the large number of generated rules.
@@
 The underlying concept of the method is a combination of linear regression and predicted models that are based on decision trees.
-This model was developed in 2005 and was positioned as a sparse linear regression model (a class of models that after the training can have coefficients with zero weights).
+This model was developed by Friedman and Popescu (2008) and was positioned as a sparse linear regression model (a class of models that after the training can have coefficients with zero weights).
 This algorithm detects the effects of interaction between features in the form of decision rules.
@@
 These rules describe the interplay of variables.
-The distinctive characteristic of LASSO is going to zero of some factors as a result of minimization.
+The distinctive characteristic of LASSO is forcing to zero the weight of some factors as a result of minimization.
 This procedure determines the most informative factors.
 The tendency of some factors to disappear is controlled with a regularization parameter that is fitted using cross-validation (procedure of empiric estimation of the generalization ability of the algorithm).
-More information about adjusting the regularization parameter can be found here (Friedman et al., 2010).
+More information about adjusting the regularization parameter can be found in Friedman et al.
+(2010).
 Due to the fact that the algorithm automatically selects the most significant factors the additional checking of selected factors is not required.
 Unlike the Ordinary Least Squares (OLS) method, the LASSO (and consequently RuleFit) does not have any prerequisites.
-We can estimate the effectiveness of the model by using common evaluation metrics such as (Mean Squared Error), (Mean Absolute Error), (Average approximation error), and (Coefficient of determination).
+We can estimate the effectiveness of the model by using common evaluation metrics such as , , , and (Mean Squared Error):
 For building the RuleFit model we used both binary and continuous features.
-As continuous features, we used the factors from the correlation matrix (Figure 10) where correlation with the dependent variable is 0,15 and more.
+As continuous features, we used the factors from the correlation matrix (Figure 10) where correlation with the dependent variable is 0.15 and more.
 Factors satisfying this are:
@@
 : Aggregate: Year built & Number of storeys
-As binary features we used the factors describing the storey of the flat (“the first”, “the last”); the wall material of the building (“bricky”, “large-panel”, “monolithic”, “small-element concrete blocks”; at the same time the wall materials such as “made of timber” and “not determined” were excluded from consideration because they are nonrepresentational and don’t have economic substance).
+As binary features we used the factors describing the storey of the flat (“the first”, “the last”); the wall material of the building (“bricky”, “large-panel”, “monolithic”, “small-element concrete blocks”; at the same time the wall materials such as “made of timber” and “not determined” were excluded from consideration because they are nonrepresentational and do not have economic substance).
 Thus we used 17 features for modeling.
@@
 In order to keep the interpretability of the model we applied a limit on the number of rules as a hyperparameter of the model – 50 rules.
-For a generation of rules we used the Random Forest algorithm.
-As a result of LASSO the part of rules were evaluated as insignificant and excluded, and only significant were kept.
+For generation of rules we used the Random Forest algorithm.
+As a result of LASSO, part of rules were evaluated as insignificant and excluded, and only the significant were kept.
 The final model contains 22 rules and 12 features.
@@
 Thus the RuleFit is a badly reproduced model and because of it, we evaluate the final results of the model on the whole set of data without splitting it on test and train.
-But in the process of study we did such splitting and made crossvalidation which gives the average to be 0,58.
+But in the process of study we did such splitting and made crossvalidation which gives the average to be 0.58.
 We present the result of the RuleFit model in Table 7.
-Average approximation error | ||
-0,6 | 13 518,28 | 8,8% |
-From Table 5 we can see that the coefficient of determination is equal to 0,6 which is acceptable for our purposes.
+0.6 | 13 518.28 | 8.8% |
+From Table 7 we can see that the coefficient of determination is equal to 0.6 which is acceptable for our purposes.
 Values of MAE and average approximation error of RuleFit are lower than for linear regression.
@@
 The successful combination of different methods in one model indicates a good perspective of further unification of well-accepted models in order to improve the efficiency of an interpretable model on complex data.
+Acknowledgments
+Authors thank Artyom S. Tanashkin for making the valuable comments and suggestions.
+Maps are drawn using OpenStreetMap (OpenStreetMap contributors, 2017) and Python library Folium.
+For data analysis and modeling we used the following Python libraries (listed in the alphabet order): GeoPy, Matplotlib (Hunter, 2007), NetworkX, NumPy (Harris et al., 2020), OSMnx, pandas (Wes McKinney, 2010), RuleFit (RuleFit contributors, ), SciKit-learn, SciKit GStat (Mälicke, 2021), SciPy, seaborn (Waskom, 2021), statsmodels.
 Appendix A The resulting model for RuleFit
@@
 : Distance to the seacoast (straight-line distance)
-: Distance to the business activities center (straight-line distance)
+: Distance to the closest business activities center (straight-line distance)
 : Distance to the closest railway station (straight-line distance)
@@
 № | Factor/Rule | Type | Coefficient |
-1 | linear | 6713,23 | |
-2 | linear | -438,15 | |
-3 | linear | -3,14 | |
-4 | linear | 6482,54 | |
-5 | linear | -2,83 | |
-6 | linear | 7435,15 | |
-7 | linear | 2,02 | |
-8 | linear | 76,72 | |
-9 | linear | -0,24 | |
-10 | linear | -2771,67 | |
-11 | linear | -387,96 | |
-12 | linear | -0,08 | |
-13 | rule | -26328,13 | |
-14 | rule | 29815,42 | |
-15 | rule | -6277,76 | |
-16 | rule | 17615,75 | |
-17 | rule | 13697,61 | |
-18 | rule | -6846,8 | |
-19 | rule | 8955,08 | |
-20 | rule | -4897,04 | |
-21 | rule | -4758,11 | |
-22 | rule | -7064,57 | |
-23 | rule | -3785,21 | |
-24 | rule | 4401,15 | |
-25 | rule | 6457,31 | |
-26 | rule | -3693,24 | |
-27 | rule | 2686,18 | |
-28 | rule | 5532,75 | |
-29 | rule | 4444,61 | |
-30 | rule | -898,74 | |
-31 | rule | 718,14 | |
-32 | rule | -387,44 | |
-33 | rule | 399,47 | |
-34 | rule | 16,4 |
+1 | linear | 6713.23 | |
+2 | linear | -438.15 | |
+3 | linear | -3.14 | |
+4 | linear | 6482.54 | |
+5 | linear | -2.83 | |
+6 | linear | 7435.15 | |
+7 | linear | 2.02 | |
+8 | linear | 76.72 | |
+9 | linear | -0.24 | |
+10 | linear | -2771.67 | |
+11 | linear | -387.96 | |
+12 | linear | -0.08 | |
+13 | rule | -26328.13 | |
+14 | rule | 29815.42 | |
+15 | rule | -6277.76 | |
+16 | rule | 17615.75 | |
+17 | rule | 13697.61 | |
+18 | rule | -6846.8 | |
+19 | rule | 8955.08 | |
+20 | rule | -4897.04 | |
+21 | rule | -4758.11 | |
+22 | rule | -7064.57 | |
+23 | rule | -3785.21 | |
+24 | rule | 4401.15 | |
+25 | rule | 6457.31 | |
+26 | rule | -3693.24 | |
+27 | rule | 2686.18 | |
+28 | rule | 5532.75 | |
+29 | rule | 4444.61 | |
+30 | rule | -898.74 | |
+31 | rule | 718.14 | |
+32 | rule | -387.44 | |
+33 | rule | 399.47 | |
+34 | rule | 16.4 |